# A Non-Abelian Approach to Riemann Surfaces

### 04 Mar 2026 v1

@@ 11.2. Quantum modular connections
If AA is a quantum modular connection and $f\in M_{0}(\Gamma,V)$ is $\Gamma$-invariant, then $A(f)$ transforms by the inhomogeneous rule
@@ 11.3. Classical modular connections
If $a_{1}$ and $a_{2}$ are classical modular connections of the same eccentricity $e$, then their difference is a weight 22 matrix modular form:
@@ 11.5. The Serre derivative
is the (holomorphic) Eisenstein series of weight 22. Unlike $E_{2m}$ for $m\geq 2$, the function $E_{2}$ is not a modular form. Instead it satisfies the quasi-modular transformation law
@@ 11.6. Ramanujan identities
defines a well-defined modular 22-differential on $X$. Indeed, the anomaly term involves the Schwarzian $S(\lambda)$ of the transition function, and $S(\lambda)=0$ for Möbius changes of coordinates. If AA is a classical modular connection with values in $\operatorname{End}(V)$, then $F_{A}\in M_{4}(\Gamma,\operatorname{End}(V))$.
@@ 11.7. Modular Curvature
Then AA is a classical modular connection of eccentricity $e=1$. Consequently its curvature
@@
is a meromorphic modular form of weight 44. Thus we obtain a natural nonlinear operator
@@ 11.8. The Chazy equation
which should be viewed as a curvature-type 22-differential. However, as we saw in Section 3, $F_{A}$ is not globally tensorial on an arbitrary Riemann surface: under changes of coordinate it acquires an anomaly term proportional to the Schwarzian derivative of the transition map. Thus one should not expect a literal analogue of the classical Bianchi identity in general. Chazy introduced in [1] a class of nonlinear third-order differential equations whose general solutions have fixed critical points. A distinguished example is the Chazy equation
@@
so that AA is a scalar classical modular connection of eccentricity $e=1$. Let $F_{A}=A^{\prime}-\frac{1}{2}A^{2}$ be the curvature. Then AA satisfies the Chazy equation
@@
In this section we apply the ideas developed so far to study periods of holomorphic 11-forms on a one-parameter family of compact Riemann surfaces of fixed genus $g$, generalizing the formula of Dedekind mentioned in the introduction. Differential equations satisfied by periods of holomorphic 11-forms often fall under the title of Picard–Fuchs equations: these are scalar linear ODEs of order $2g$ obtained from the Gauss–Manin system. While extremely useful, such scalar equations are not well suited for the Schwarzian point of view pursued in this paper. For $g>1$ we therefore propose a different approach which produces a canonical second order ODE with $g\times g$ matrix coefficients in $\operatorname{End}(E)$, and from there we obtain new insight about periods using the matrix Schwarzian derivative. Such a one-parameter family, roughly speaking, either corresponds to a non-constant holomorphic map from a Riemann surface $S$ to the moduli space of curves of genus $g$, $\mathcal{M}_{g}$, or equivalently to a holomorphic fibration $\pi\colon X\to S$ from a complex surface $X$ to $S$ whose fibers $X_{s}=\pi^{-1}(s)$ are Riemann surfaces of genus $g$.
@@
where $H^{1,0}(C)\simeq H^{0}(C,K_{C})$ is the space of holomorphic 11-forms and $H^{0,1}(C)\simeq\overline{H^{1,0}(C)}$ is the space of anti-holomorphic 11-forms.
@@
is the Hodge bundle whose fiber at $s\in S$ is $H^{1,0}(X_{s})$. We will also construct a classical 22-differential
@@
will then encode the variation of periods: for a locally constant homology class $\gamma\in H_{1}(X_{s},\mathbb{Z})$ and a holomorphic 11-form $\omega_{s}\in H^{1,0}(X_{s})$ varying holomorphically with $s$, the period function
@@ 12. Higher genus periods
Lemma 12.4.
In the setting of Definition 12.3, the collections $\{A_{U}\}$ and $\{q_{U}\}$ glue to define a classical connection $A=A_{\phi}$ on $E$ of eccentricity $e=\tfrac{1}{2}$ and a classical 22-differential $q=q_{\phi}$ with values in $\operatorname{End}(E)$.
@@
In particular, AA transforms as a (classical) connection of eccentricity $e$ and $q$ transforms as an $\operatorname{End}(E)$-valued 22-differential.
@@
Fix a point of $U_{i}\cap U_{j}$ and write $\lambda=\lambda_{ij}$. Set $p_{i}:=2A_{i}$. If we represent the same section in the $(U_{j},s_{j})$-chart by
$\widetilde{\omega}:=\omega^{(i)}\circ\lambda,$
then $\widetilde{\omega}$ satisfies a second-order equation of the form
@@
with respect to the coordinate $s_{j}$. By Lemma 7.1 (applied componentwise to the vector–valued function $\widetilde{\omega}$), the coefficients are
@@
Writing $\widetilde{p}_{j}=2\widetilde{A}_{j}$, we obtain
@@
Finally, the chosen frame on $U_{j}$ is
$\omega^{(j)}=g_{ji}\,\widetilde{\omega},\qquad g_{ji}=g_{ij}^{-1},$
with $g_{ji}$ constant. Hence differentiation commutes with multiplication by $g_{ji}$, and $\omega^{(j)}$ satisfies
@@
Substituting the expressions for $\widetilde{A}_{j}$ and $\widetilde{q}_{j}$ yields the stated transformation laws. ∎
@@ 13. A genus two family
We conclude the section with a concrete genus two example. Recall that every smooth curve of genus 22 is hyperelliptic, so hyperelliptic families provide a natural and sufficiently general testing ground for the matrix Schwarzian formalism.
Let $U\subset\mathbb{C}$ be a simply connected open set avoiding the discriminant locus, and consider the one–parameter family of genus 22 curves
@@
On $X_{t}$ the holomorphic 11-forms are spanned by
@@
A direct computation in $H^{1}_{\mathrm{dR}}(X_{t})$ (for instance via Griffiths reduction) shows that the 44 classes
@@
Since $e=\tfrac{1}{2}$, the curvature of AA is $F_{A}=A^{\prime}-A^{2}$ and the (matrix) quantum Schwarzian associated to (13.2) is
@@
The matrix Schwarzian $\mathcal{S}_{A,q}$ produces scalar invariants by taking characteristic combinations. In the present genus two case one obtains in particular the 22-differentials
@@ 14. Cubic threefolds in $\mathbb{P}^{4}$
In this section we illustrate how the second order non-abelian formalism developed in Section 7 and Part IV extends to a basic higher-dimensional moduli problem, namely smooth cubic threefolds in $\mathbb{P}^{4}$. While the Gauss–Manin connection for a one-parameter family of cubic threefolds is a first-order system of rank $10$ on $H^{3}$, our approach produces a canonical second order system on the rank-55 Hodge bundle $E=H^{2,1}$ with coefficients in $\operatorname{End}(E)$. For a natural deformation of the Fermat cubic, a large symmetry group forces this system to split into one-dimensional character spaces; in a symmetry-adapted basis the coefficient matrices become diagonal. We explain the geometric input (Hodge theory, residues, and the Gauss–Manin system), state the resulting diagonal second order ODE, and omit only the standard reduction computations.
@@ 14.1. Cubic threefolds
A cubic threefold is a smooth hypersurface $X\subset\mathbb{P}^{4}$ defined by a homogeneous polynomial $F(x_{0},\dots,x_{4})=0$ of degree 33. Two classical examples are the Fermat cubic threefold
@@
See for instance [3, 20].
Since $h^{3,0}=0$, the Hodge structure on $H^{3}(X)$ is of type $(2,1)+(1,2)$. Writing $M:=X$, the associated intermediate Jacobian can be presented as the complex torus
$J(M):=H^{1,2}(M)\big/H^{3}(M,\mathbb{Z}),$
where we view $H^{3}(M,\mathbb{Z})$ as a lattice in $H^{1,2}(M)$ via the projection $H^{3}(M,\mathbb{C})\twoheadrightarrow H^{1,2}(M)$ coming from the Hodge decomposition $H^{3}(M,\mathbb{C})=H^{2,1}(M)\oplus H^{1,2}(M)$. It is a principally polarized abelian variety of dimension 55 [3]. The intermediate Jacobian plays a central role in the geometry of cubic threefolds (for instance in the Clemens–Griffiths irrationality theorem).
@@ 14.2. Fermat cubic
Let $\Omega$ be the standard homogeneous 44-form on $\mathbb{C}^{5}$,
@@ 14.3. Canonical second-order equation
For $i=0,1,2$ the corresponding degree–44 numerator lies in the Jacobian ideal of $F_{t}$, hence its residue class vanishes; this yields $\omega_{0}^{\prime}=\omega_{1}^{\prime}=\omega_{2}^{\prime}=0$. For $i=3,4$ one differentiates once more, obtaining a residue with pole order 44 whose degree–77 numerator is a multiple of $(x_{0}x_{1}x_{2})^{2}$. Using the relations $\partial_{x_{0}}F_{t}=3(x_{0}^{2}-tx_{1}x_{2})$, $\partial_{x_{1}}F_{t}=3(x_{1}^{2}-tx_{0}x_{2})$, $\partial_{x_{2}}F_{t}=3(x_{2}^{2}-tx_{0}x_{1})$, together with the Griffiths–Dwork reduction identity [4, 7], one rewrites this pole–44 class in the span of $\omega_{i}$ and $\omega_{i}^{\prime}$. In the present monomial deformation, the reduction is particularly simple and produces the diagonal coefficients (14.5). ∎
@@
15. The mass–spring system
@@ 15.1. What is time?
15.2. Classical mass–spring systems
We begin with a frictionless unit mass attached to a spring of stiffness $k>0$. In the usual sign convention the displacement $\psi$ satisfies $\ddot{\psi}+k\psi=0$. To match the normalization used throughout the paper, we rewrite this as
@@ 15.2. Classical mass–spring systems
Now imagine that $p$ and $q$ vary meromorphically with respect to (complex) time. In a local clock $t$ we then have
@@ 15.3. Eternal mass–spring system
where II denotes the identity matrix. In particular, $p$ transforms affinely (with an inhomogeneous $\lambda^{\prime\prime}/\lambda^{\prime}$ term), while $q$ transforms tensorially with weight 22. In our notation, for an eternal system the quantity $A=p/2$ transforms from one chart to another as a time-connection with eccentricity $e=\tfrac{1}{2}$, while $q$ transforms as a time-quadratic differential.
A natural coordinate-free invariant is obtained from the Schwarzian curvature. For $e=\tfrac{1}{2}$ the curvature of AA is

### 14 Mar 2026 v2

@@ 11.2. Quantum modular connections
If $A$ is a quantum modular connection and $f\in M_{0}(\Gamma,V)$ is $\Gamma$-invariant, then $A(f)$ transforms by the inhomogeneous rule
@@ 11.3. Classical modular connections
If $a_{1}$ and $a_{2}$ are classical modular connections of the same eccentricity $e$, then their difference is a weight $2$ matrix modular form:
@@ 11.5. The Serre derivative
is the (holomorphic) Eisenstein series of weight $2$. Unlike $E_{2m}$ for $m\geq 2$, the function $E_{2}$ is not a modular form. Instead it satisfies the quasi-modular transformation law
@@ 11.6. Ramanujan identities
defines a well-defined modular $2$-differential on $X$. Indeed, the anomaly term involves the Schwarzian $S(\lambda)$ of the transition function, and $S(\lambda)=0$ for Möbius changes of coordinates. If $A$ is a classical modular connection with values in $\operatorname{End}(V)$, then $F_{A}\in M_{4}(\Gamma,\operatorname{End}(V))$.
@@ 11.7. Modular Curvature
Then $A$ is a classical modular connection of eccentricity $e=1$. Consequently its curvature
@@
is a meromorphic modular form of weight $4$. Thus we obtain a natural nonlinear operator
@@ 11.8. The Chazy equation
which should be viewed as a curvature-type $2$-differential. However, as we saw in Section 3, $F_{A}$ is not globally tensorial on an arbitrary Riemann surface: under changes of coordinate it acquires an anomaly term proportional to the Schwarzian derivative of the transition map. Thus one should not expect a literal analogue of the classical Bianchi identity in general. Chazy introduced in [1] a class of nonlinear third-order differential equations whose general solutions have fixed critical points. A distinguished example is the Chazy equation
@@
so that $A$ is a scalar classical modular connection of eccentricity $e=1$. Let $F_{A}=A^{\prime}-\frac{1}{2}A^{2}$ be the curvature. Then $A$ satisfies the Chazy equation
@@
In this section we apply the ideas developed so far to study periods of holomorphic $1$-forms on a one-parameter family of compact Riemann surfaces of fixed genus $g$, generalizing the formula of Dedekind mentioned in the introduction. Differential equations satisfied by periods of holomorphic $1$-forms often fall under the title of Picard–Fuchs equations: these are scalar linear ODEs of order $2g$ obtained from the Gauss–Manin system. While extremely useful, such scalar equations are not well suited for the Schwarzian point of view pursued in this paper. For $g>1$ we therefore propose a different approach which produces a canonical second order ODE with $g\times g$ matrix coefficients in $\operatorname{End}(E)$, and from there we obtain new insight about periods using the matrix Schwarzian derivative. Such a one-parameter family, roughly speaking, either corresponds to a non-constant holomorphic map from a Riemann surface $S$ to the moduli space of curves of genus $g$, $\mathcal{M}_{g}$, or equivalently to a holomorphic fibration $\pi\colon X\to S$ from a complex surface $X$ to $S$ whose fibers $X_{s}=\pi^{-1}(s)$ are Riemann surfaces of genus $g$.
@@
where $H^{1,0}(C)\simeq H^{0}(C,K_{C})$ is the space of holomorphic $1$-forms and $H^{0,1}(C)\simeq\overline{H^{1,0}(C)}$ is the space of anti-holomorphic $1$-forms.
@@
is the Hodge bundle whose fiber at $s\in S$ is $H^{1,0}(X_{s})$. We will also construct a classical $2$-differential
@@
will then encode the variation of periods: for a locally constant homology class $\gamma\in H_{1}(X_{s},\mathbb{Z})$ and a holomorphic $1$-form $\omega_{s}\in H^{1,0}(X_{s})$ varying holomorphically with $s$, the period function
@@ 12. Higher genus periods
Theorem 12.4.
In the setting of Definition 12.3, the collections $\{A_{U}\}$ and $\{q_{U}\}$ glue to define a classical connection $A=A_{\phi}$ on $E$ of eccentricity $e=1/2$ and a classical $2$-differential $q=q_{\phi}$ with values in $\operatorname{End}(E)$.
@@
In particular, $A$ transforms as a (classical) connection of eccentricity $e$ and $q$ transforms as an $\operatorname{End}(E)$-valued $2$-differential.
@@
The key point is that the relative de Rham bundle
$\mathcal{H}=H^{1}_{\mathrm{dR}}(X/S)$
carries the flat Gauss–Manin connection. Hence on each simply connected chart $U_{i}$ we may choose a flat trivialization of $\mathcal{H}|_{U_{i}}$, and on an overlap $U_{i}\cap U_{j}$ two such trivializations differ by a locally constant matrix. After shrinking if necessary, we may therefore write the two Hodge frames in the form
$\omega^{(j)}=g_{ji}\,(\omega^{(i)}\circ\lambda_{ij}),\qquad g_{ji}=g_{ij}^{-1},$
with $g_{ij}\in\operatorname{GL}_{g}(\mathbb{C})$ locally constant on $U_{i}\cap U_{j}$. This justifies the constant transition matrix appearing in the statement.
We next verify that the second-order system is canonical. On a chart $(U_{i},s_{i})$, genericity implies that
$\omega^{(i)}_{1},\dots,\omega^{(i)}_{g},\;\partial_{s_{i}}\omega^{(i)}_{1},\dots,\partial_{s_{i}}\omega^{(i)}_{g}$
form a frame of $\mathcal{H}|_{U_{i}}$ on a dense open subset. Therefore $(\omega^{(i)},(\omega^{(i)})^{\prime})$ is a basis there, and so the second derivative $(\omega^{(i)})^{\prime\prime}$ can be written uniquely as
$(\omega^{(i)})^{\prime\prime}=2A_{i}\,(\omega^{(i)})^{\prime}+q_{i}\,\omega^{(i)}.$
Thus the matrices $A_{i}$ and $q_{i}$ are uniquely determined by the family and the chosen coordinate, so the resulting second-order system is canonical once the Hodge frame is fixed.
Now fix a point of $U_{i}\cap U_{j}$ and write $\lambda=\lambda_{ij}=s_{i}\circ s_{j}^{-1}$. Set $p_{i}:=2A_{i}$. Consider first the same vector-valued section written in the $s_{j}$-coordinate but before changing the flat frame:
$\widetilde{\omega}:=\omega^{(i)}\circ\lambda.$
By Lemma 7.1, applied componentwise to the vector-valued function $\widetilde{\omega}$, the equation
$(\omega^{(i)})^{\prime\prime}=p_{i}\,(\omega^{(i)})^{\prime}+q_{i}\,\omega^{(i)}$
transforms into
@@
with
@@
Writing $\widetilde{p}_{j}=2\widetilde{A}_{j}$, we get
@@
Finally, passing from $\widetilde{\omega}$ to the chosen Hodge frame
$\omega^{(j)}=g_{ji}\,\widetilde{\omega}$,
with $g_{ji}$ locally constant, differentiation commutes with multiplication by $g_{ji}$. Hence $\omega^{(j)}$ satisfies
@@
Substituting the formulas for $\widetilde{A}_{j}$ and $\widetilde{q}_{j}$ gives
$A_{j}=g_{ji}\Bigl(\lambda^{\prime}(A_{i}\circ\lambda)+\frac{1}{2}\frac{\lambda^{\prime\prime}}{\lambda^{\prime}}I\Bigr)g_{ij},\qquad q_{j}=(\lambda^{\prime})^{2}\,g_{ji}\,(q_{i}\circ\lambda)\,g_{ij},$
which is exactly the claimed transformation law. Therefore the local matrices $\{A_{U}\}$ glue to a classical connection of eccentricity $e=\tfrac{1}{2}$ on $E$, and the local matrices $\{q_{U}\}$ glue to an $\operatorname{End}(E)$-valued classical $2$-differential. ∎
@@ 13. A genus two family
We conclude the section with a concrete genus two example. Recall that every smooth curve of genus $2$ is hyperelliptic, so hyperelliptic families provide a natural and sufficiently general testing ground for the matrix Schwarzian formalism.
Let $U\subset\mathbb{C}$ be a simply connected open set avoiding the discriminant locus, and consider the one–parameter family of genus $2$ curves
@@
On $X_{t}$ the holomorphic $1$-forms are spanned by
@@
A direct computation in $H^{1}_{\mathrm{dR}}(X_{t})$ (for instance via Griffiths reduction) shows that the $4$ classes
@@
Since $e=\tfrac{1}{2}$, the curvature of $A$ is $F_{A}=A^{\prime}-A^{2}$ and the (matrix) quantum Schwarzian associated to (13.2) is
@@
The matrix Schwarzian $\mathcal{S}_{A,q}$ produces scalar invariants by taking characteristic combinations. In the present genus two case one obtains in particular the $2$-differentials
@@ 14. Cubic threefolds in $\mathbb{P}^{4}$
In this section we illustrate how the second order non-abelian formalism developed in Section 7 and Part IV extends to a basic higher-dimensional moduli problem, namely smooth cubic threefolds in $\mathbb{P}^{4}$. While the Gauss–Manin connection for a one-parameter family of cubic threefolds is a first-order system of rank $10$ on $H^{3}$, our approach produces a canonical second order system on the rank-$5$ Hodge bundle $E=H^{2,1}$ with coefficients in $\operatorname{End}(E)$. For a natural deformation of the Fermat cubic, a large symmetry group forces this system to split into one-dimensional character spaces; in a symmetry-adapted basis the coefficient matrices become diagonal. We explain the geometric input (Hodge theory, residues, and the Gauss–Manin system), state the resulting diagonal second order ODE, and omit only the standard reduction computations.
@@ 14.1. Cubic threefolds
A cubic threefold is a smooth hypersurface $X\subset\mathbb{P}^{4}$ defined by a homogeneous polynomial $F(x_{0},\dots,x_{4})=0$ of degree $3$. Two classical examples are the Fermat cubic threefold
@@
Equivalently, the Hodge diamond has the form shown in Figure 1; see for instance [3, 22].
1100011005555001100011
Figure 1. Hodge diamond of a smooth cubic threefold.
Since $h^{3,0}=0$, the Hodge structure on $H^{3}(X)$ is of type $(2,1)+(1,2)$. The associated intermediate Jacobian can be presented as the complex torus
$J(X):=H^{1,2}(X)\big/H^{3}(X,\mathbb{Z}),$
where we view $H^{3}(X,\mathbb{Z})$ as a lattice in $H^{1,2}(X)$ via the projection $H^{3}(X,\mathbb{C})\twoheadrightarrow H^{1,2}(X)$ coming from the Hodge decomposition $H^{3}(X,\mathbb{C})=H^{2,1}(X)\oplus H^{1,2}(X)$. It is a principally polarized abelian variety of dimension $5$ [3]. The intermediate Jacobian plays a central role in the geometry of cubic threefolds (for instance in the Clemens–Griffiths irrationality theorem).
@@ 14.2. Fermat cubic
This family is chosen because it gives a fully explicit computable model of the canonical second-order system.
@@
Let $\Omega$ be the standard homogeneous $4$-form on $\mathbb{C}^{5}$,
@@ 14.3. Canonical second-order equation
For $i=0,1,2$ the corresponding degree–$4$ numerator lies in the Jacobian ideal of $F_{t}$, hence its residue class vanishes; this yields $\omega_{0}^{\prime}=\omega_{1}^{\prime}=\omega_{2}^{\prime}=0$. For $i=3,4$ one differentiates once more, obtaining a residue with pole order $4$ whose degree–$7$ numerator is a multiple of $(x_{0}x_{1}x_{2})^{2}$. Using the relations $\partial_{x_{0}}F_{t}=3(x_{0}^{2}-tx_{1}x_{2})$, $\partial_{x_{1}}F_{t}=3(x_{1}^{2}-tx_{0}x_{2})$, $\partial_{x_{2}}F_{t}=3(x_{2}^{2}-tx_{0}x_{1})$, together with the Griffiths–Dwork reduction identity [4, 7], one rewrites the cohomology class represented by this rational form with pole order $4$ in the span of $\omega_{i}$ and $\omega_{i}^{\prime}$. In the present monomial deformation, the reduction is particularly simple and produces the diagonal coefficients (14.5). ∎
@@
15. The classical mass–spring system
@@ 15.1. What is time?
15.2. The Hooke’s law
We begin with a frictionless unit mass attached to a spring of stiffness $k>0$. The restoring force is given by Hooke’s law:
(15.1) $F=-k\,\psi.$
For a unit mass, Newton’s equation therefore yields
$\ddot{\psi}+k\psi=0.$
To match the normalization used throughout the paper, we rewrite this as
@@ 15.2. The Hooke’s law
Now imagine that $p$ and $q$ vary meromorphically with respect to time. In a local clock $t$ we then have
@@ 15.3. Eternal mass–spring system
where $I$ denotes the identity matrix. In particular, $p$ transforms affinely (with an inhomogeneous $\lambda^{\prime\prime}/\lambda^{\prime}$ term), while $q$ transforms tensorially with weight $2$. In our notation, for an eternal system the quantity $A=p/2$ transforms from one chart to another as a time-connection with eccentricity $e=\tfrac{1}{2}$, while $q$ transforms as a time-quadratic differential. A natural coordinate-free invariant is obtained from the Schwarzian curvature. For $e=\tfrac{1}{2}$ the curvature of $A$ is